# Near-Field Channel Estimation and Joint Angle-Range Recovery in XL-MIMO Systems: A Gridless Super-Resolution Approach

### v1

@@ V-B Performance under Different SNR Levels
We now evaluate channel, angle, and range estimation performance across various SNR levels and with different range configurations. As shown in Figs. 3(a) and 3(d), the NMSE of ANM(SI) decreases monotonically with increasing SNR, whereas the on-grid SW-OMP and P-SOMP algorithms exhibit error floors at moderate-to-high SNR due to basis mismatch. For SNR >> 5 dB, ANM(SI) significantly outperforms the on-grid methods, demonstrating the advantage of the proposed gridless super-resolution approach. By incorporating the refinement step in Stage-II of Algorithm 1, ANM(SI+SII) achieves lower NMSE than ANM(SI) and substantially outperforms the off-grid SS-SIGW-OLS and P-SIGW algorithms when SNR >> 15 dB. In the low SNR regime (5-5 to $5$ dB), ANM(SI) performs slightly worse than the polar-domain P-SOMP, and ANM(SI+SII) is marginally inferior to P-SIGW when SNR <15<15 dB. As indicated by Figs. 3(b) and 3(e), this behavior likely stems from less accurate Stage-I angle estimates at low SNR, which limits the effectiveness of Stage-II refinement.
@@ V-C Performance with Different Range Parameters
We now investigate the impact of the distance between the scatterers and the BS on the performance of the proposed near-field channel estimation method. The distance varies from 5m to 120m, covering the near-field region and extending toward the far-field boundary. In the simulations, we set SNR = $30$ dB. Fig. 4(a) shows the NMSE of channel estimation versus distance. The NMSE of ANM(SI) remains nearly constant at approximately 60-60 dB across the entire range, indicating that the Stage-I reconstruction is largely insensitive to user location. In contrast, ANM(SI+SII) achieves substantially lower NMSE once the distance exceeds 20m, demonstrating the effectiveness of the Stage-II refinement when the user is not extremely close to the BS. The on-grid P-SOMP suffers from a persistent high error floor due to basis mismatch, whereas SW-OMP and SS-SIGW-OLS improve with increasing distance but remain ineffective in the deep near-field region.
@@ V-D Performance with Different Pilot Lengths
We next evaluate the impact of pilot length $Q$ on the performance of the proposed near-field channel estimation method. The pilot length varies from $Q=16$ to $Q=64$, corresponding to different training overheads. We fix the SNR at 30 dB. As shown in Fig. 5(a), the NMSEs of both ANM(SI) and ANM(SI+SII) improve steadily as $Q$ increases, and the curves gradually saturate for larger $Q$. These results demonstrate that the proposed methods effectively exploit additional pilot observations. In contrast, the on-grid P-SOMP and the off-grid P-SIGW exhibit only marginal improvement with increasing pilot lengths, whereas SW-OMP and SS-SIGW-OLS maintain nearly constant and significantly higher NMSE across all values of $Q$. The angle and range estimation results in Figs. 5(b) and 5(c) follow a similar trend. Both ANM(SI) and ANM(SI+SII) benefit substantially from longer pilot sequences, further confirming that the proposed algorithms capture the spherical wavefront structure more accurately as more observations become available.
@@ VI Conclusion
$0\in\partial\|\widehat{\mathbf{X}}\|_{\mathcal{A}}+\frac{1}{\tau}\mathcal{P}^{\ast}_{y}(\mathcal{P}(\widehat{\mathbf{X}})-\mathbf{y}).$ (-B.1)
There exists a dual variable $\mathbf{S}\in\partial\|\widehat{\mathbf{X}}\|_{\mathcal{A}}$ with $\|\mathbf{S}\|^{\ast}_{\mathcal{A}}\leq 1$ and $\langle\mathbf{S},\widehat{\mathbf{X}}\rangle_{\mathbb{R}}=\|\widehat{\mathbf{X}}\|_{\mathcal{A}}$. Thus, we have $\mathbf{S}=-\frac{1}{\tau}\mathcal{P}^{\ast}_{y}(\mathcal{P}(\widehat{\mathbf{X}})-\mathbf{y})$.
@@
[…]
-C Proof of Proposition 2
Let $\mathbf{Z}=\mathcal{P}_{y}^{*}(\mathbf{n})\in\mathbb{C}^{P\times N}$ and define $\mathbf{X}(\omega)=\mathbf{Z}\mathbf{a}(\omega)\in\mathbb{C}^{P}$. For a unit norm $\mathbf{u}\in\mathbb{C}^{P}$, i.e., $\|\mathbf{u}\|_{2}=1$, define the scalar process
$Z(\omega,\mathbf{u})=\langle\mathbf{u},\mathbf{X}(\omega)\rangle_{\mathbb{R}}=\Big\langle\mathbf{n},\ \mathcal{P}_{y}\!\big(\mathbf{u}\mathbf{a}^{H}(\omega)\big)\Big\rangle_{\mathbb{R}}.$ (-C.1)
Since $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M})$ and $\|\mathcal{P}_{y}(\mathbf{u}\mathbf{a}^{H}(\omega))\|_{2}\leq\|\mathcal{P}_{y}\|_{F\to 2}\,\|\mathbf{u}\mathbf{a}^{H}(\omega)\|_{F}\leq\sqrt{N}\,\|\mathcal{P}_{y}\|_{F\to 2}$, each $Z(\omega,\mathbf{u})$ is zero-mean sub-Gaussian with parameter at most $c_{0}\,\sigma\,\sqrt{N}\,\|\mathcal{P}_{y}\|_{F\to 2}$, where $c_{0}$ is a constant. To prove the results in Proposition 2, we have to derive the bound of $\Big\|\mathcal{P}_{y}^{*}(\mathbf{n})\Big\|_{\mathcal{A}}^{*}=\sup_{\omega\in[0,1)}\|\mathbf{X}(\omega)\|_{2}$. To proceed, we bound the original supremum by first restricting it to a finite grid of points, then applying a $(1/2)$-net on the unit sphere in $\mathbb{C}^{P}$ using Lemmas -C.1 and -C.2, and finally controlling the finite maximum via sub-Gaussian tail bounds.
Lemma -C.1.
Let $\mathbb{S}^{P-1}=\{\mathbf{u}\in\mathbb{C}^{P}:\|\mathbf{u}\|_{2}=1\}$ be the unit sphere and $\mathcal{U}\subset\mathbb{S}^{P-1}$ be a $(1/2)$-net of $\mathbb{S}^{P-1}$. Then, there exists a $(1/2)$-net $\mathcal{U}$ with cardinality
$|\mathcal{U}|\;\leq\;5^{2P}.$ (-C.2)
Furthermore, for any $\mathbf{x}\in\mathbb{C}^{P}$, we have
$\sup_{\|\mathbf{u}\|_{2}=1}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}\;\leq\;2\max_{\mathbf{u}\in\mathcal{U}}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}.$ (-C.3)
Proof:
Note that (-C.2) follows directly from Lemma 5.2 of [Vershynin2012]. Thus, we only prove (-C.3) here.
Consider an arbitrary vector $\mathbf{x}\in\mathbb{C}^{P}$ with its $\ell_{2}$-norm defined as $\|\mathbf{x}\|_{2}=\sup_{\|\mathbf{u}\|_{2}=1}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}$. Let $\mathbf{u}\in\mathbb{S}^{P-1}$ be a vector on the unit sphere. Since $\mathcal{U}$ is a $(1/2)$-net, there exists a point $\mathbf{u}_{\varepsilon}\in\mathcal{U}$ such that $\|\mathbf{u}-\mathbf{u}_{\varepsilon}\|_{2}\leq\tfrac{1}{2}$.
By writing $\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}=\langle\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}+\langle\mathbf{u}-\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}$ and applying the Cauchy–Schwarz inequality, we have $\big|\langle\mathbf{u}-\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}\big|\leq\|\mathbf{u}-\mathbf{u}_{\varepsilon}\|_{2}\,\|\mathbf{x}\|_{2}\leq\varepsilon\,\|\mathbf{x}\|_{2}.$ Thus, we have
$\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}\leq\langle\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}+\varepsilon\,\|\mathbf{x}\|_{2}.$ (-C.4)
Taking the supremum over $\mathbf{u}\in\mathbb{S}^{P-1}$ on the left-hand side of (-C.4), it yields
$\sup_{\|\mathbf{u}\|_{2}=1}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}$ $\leq\sup_{\|\mathbf{u}\|_{2}=1}\Big(\langle\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}+\varepsilon\,\|\mathbf{x}\|_{2}\Big)$
$\leq\max_{\mathbf{u}_{\varepsilon}\in\mathcal{U}}\langle\mathbf{u}_{\varepsilon},\mathbf{x}\rangle_{\mathbb{R}}+\varepsilon\,\|\mathbf{x}\|_{2}.$ (-C.5)
Thus, we have $\|\mathbf{x}\|_{2}\leq\max_{\mathbf{u}\in\mathcal{U}}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}+\varepsilon\,\|\mathbf{x}\|_{2}$. For $\varepsilon=1/2$, this becomes $\|\mathbf{x}\|_{2}\;\leq\;2\,\max_{\mathbf{u}\in\mathcal{U}}\langle\mathbf{u},\mathbf{x}\rangle_{\mathbb{R}}$, which proves (-C.3).
Lemma -C.2.
Let $\Omega_{N}=\{k/(16N):\,k=0,\ldots,16N-1\}$, where $N\geq 1$ is an integer. For each fixed $\mathbf{u}$, the mapping $\omega\mapsto Z(\omega,\mathbf{u})$ in (-C.1) is a trigonometric polynomial of degree at most N1N-1. Furthermore, we have
$\sup_{\omega}|Z(\omega,\mathbf{u})|\leq\ 2\,\max_{\omega_{k}\in\Omega_{N}}|Z(\omega_{k},\mathbf{u})|.$ (-C.6)
Proof:
Using the canonical basis $\{\mathbf{e}_{n}\}_{n=0}^{N-1}$ of $\mathbb{C}^{N}$, we have $\mathbf{a}(\omega)=\sum_{n=0}^{N-1}e^{j2\pi n\omega}\,\mathbf{e}_{n}$. Thus, $\mathbf{u}\mathbf{a}^{H}(\omega)=\sum_{n=0}^{N-1}e^{-j2\pi n\omega}\,\mathbf{u}\mathbf{e}_{n}^{H}$. By the linearity property of $\mathcal{P}_{y}$, we have
$\mathcal{P}_{y}\big(\mathbf{u}\mathbf{a}^{H}(\omega)\big)=\sum_{n=0}^{N-1}e^{-j2\pi n\omega}\,\mathcal{P}_{y}\big(\mathbf{u}\mathbf{e}_{n}^{H}\big).$
According to (-C.1), $Z(\omega,\mathbf{u})$ can be expressed as
$\begin{split}Z(\omega,\mathbf{u})&=\sum_{n=0}^{N-1}e^{-j2\pi n\omega}\,\Big\langle\mathbf{n},\,\mathcal{P}_{y}\big(\mathbf{u}\mathbf{e}_{n}^{H}\big)\Big\rangle_{\mathbb{R}}\\ &=\sum_{n=0}^{N-1}c_{n}(\mathbf{u})\,e^{-j2\pi n\omega},\end{split}$ (-C.7)
where $c_{n}(\mathbf{u})\ =\langle\mathbf{n},\,\mathcal{P}_{y}\big(\mathbf{u}\mathbf{e}_{n}^{H}\big)\rangle_{\mathbb{R}}$ denotes the coefficient of trigonometric polynomial. Therefore, $Z(\omega,\mathbf{u})$ is a trigonometric polynomial of degree at most N1N-1 in $\omega$.
Fix $\mathbf{u}$ with $\|\mathbf{u}\|_{2}=1$ and define $f(\omega)=Z(\omega,\mathbf{u})=\sum_{n=0}^{N-1}c_{n}e^{-j2\pi n\omega}$ with $\omega\in[0,1)$. For simplicity, further define $S=\sup_{\omega\in[0,1)}|f(\omega)|$ and $M=\max_{\omega_{k}\in\Omega_{N}}|f(\omega_{k})|$. Since $f^{\prime}(\omega)=\sum_{n=0}^{N-1}c_{n}(-j2\pi n)e^{-j2\pi n\omega}$, we obtain
$\sup_{\omega\in[0,1)}|f^{\prime}(\omega)|\leq\;2\pi(N-1)\,\sup_{\omega\in[0,1)}|f(\omega)|\leq 2\pi NS.$ (-C.8)
The grid points in the set $\Omega_{N}$ have a spacing $\Delta=1/(16N)$. Thus, for any $\omega\in[0,1)$, there exists a grid point $\omega_{k}\in\Omega_{N}$ satisfying $|\omega-\omega_{k}|\leq\Delta/2\leq\Delta$. Therefore, $|f(\omega)-f(\omega_{k})|$ can be bounded as follows
[…]
$\leq\Delta\,\sup_{\widetilde{\omega}}|f^{\prime}(\widetilde{\omega})|\leq\frac{\pi}{8}\,S.$ (-C.9)
Thus, for any $\omega\in[0,1)$,
$|f(\omega)|\leq|f(\omega_{k})|+\frac{\pi}{8}S\leq M+\frac{\pi}{8}S.$
Taking the supremum over $\omega$ yields $S=\sup_{\omega}|f(\omega)|\leq M+\frac{\pi}{8}S$, i.e., $S\leq\frac{1}{1-\pi/8}\,M<2M$. In particular, $\sup_{\omega}|f(\omega)|\leq 2\max_{\omega_{k}\in\Omega_{N}}|f(\omega_{k})|$. This completes the proof. ∎
Proof of Proposition 2: Now we derive the bound for $\Big\|\mathcal{P}_{y}^{*}(\mathbf{n})\Big\|_{\mathcal{A}}^{*}=\sup_{\omega\in[0,1)}\|\mathbf{X}(\omega)\|_{2}$. Applying Lemmas -C.1 and -C.2, we have
$\begin{split}\sup_{\omega}\|\mathbf{X}(\omega)\|_{2}&=\sup_{\omega}\sup_{\|\mathbf{u}\|_{2}=1}\ Z(\omega,\mathbf{u})\\ &\leq\ 2\,\max_{\omega_{k}\in\Omega_{N}}\ \sup_{\|\mathbf{u}\|_{2}=1}Z(\omega_{k},\mathbf{u})\\ &\leq\ 4\,\max_{\omega_{k}\in\Omega_{N},\ \mathbf{u}\in\mathcal{U}}Z(\omega_{k},\mathbf{u}),\end{split}$ (-C.10)
Thus, controlling the supremum on the left-hand side of (-C.10) reduces to bounding the maximum of $Z(\omega_{k},\mathbf{u})$ over the finite index set $\mathcal{I}\ =\ \Omega_{N}\times\mathcal{U}$ with $|$ $\mathcal{I}|\ \leq\ 16N\cdot 5^{2P}$.
For each fixed $(\omega_{k},\mathbf{u})$, $Z(\omega_{k},\mathbf{u})$ is sub-Gaussian with parameter $\kappa\leq c_{0}\,\sigma\,\sqrt{N}\,\|\mathcal{P}_{y}\|_{F\to 2}$. Setting $t=\kappa\Big(\sqrt{2\log|\mathcal{I}|}+\sqrt{2\log(1/\delta)}\Big)$, we have $\frac{t^{2}}{2\kappa^{2}}\geq\log|\mathcal{I}|+\log(1/\delta)$. Applying the standard results of bounding the maximum of sub-Gaussian random variables from [Rigollet2023, Theorem 1.14], we have
$\mathbb{E}\,\sup_{\omega}\|\mathbf{X}(\omega)\|_{2}\ \leq\ C\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,q_{N,P},$ (-C.11)
$\begin{split}&\mathbb{P}\!\left\{\sup_{\omega}\|\mathbf{X}(\omega)\|_{2}\leq C^{\prime}\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,\big(q_{N,P}+\sqrt{N\log(1/\delta)}\big)\right\}\\ &\geq\ 1-\delta,\end{split}$ (-C.12)
where $C,C^{\prime}$ are constants.
Recalling $\|\mathcal{P}_{y}^{*}(\mathbf{n})\|_{\mathcal{A}}^{*}=\sup_{\omega}\|\mathbf{X}(\omega)\|_{2}$ proves (25) and (26). $\square$

### v2

@@ V-B Performance under Different SNR Levels
We now evaluate channel, angle and range estimation performance across various SNR levels and with different range configurations. As shown in Figs. 3(a) and 3(d), the NMSE of ANM(SI) decreases monotonically with increasing SNR, whereas the on-grid SW-OMP and P-SOMP algorithms exhibit error floors at moderate-to-high SNR due to basis mismatch. For SNR >> 5 dB, ANM(SI) significantly outperforms the on-grid methods, demonstrating the advantage of the proposed gridless super-resolution approach. By incorporating the refinement step in Stage-II of Algorithm 1, ANM(SI+SII) achieves lower NMSE than ANM(SI) and substantially outperforms the off-grid SS-SIGW-OLS and P-SIGW algorithms when SNR >> 15 dB. In the low SNR regime (5-5 to $5$ dB), ANM(SI) performs slightly worse than the polar-domain P-SOMP, and ANM(SI+SII) is marginally inferior to P-SIGW when SNR <15<15 dB. As indicated by Figs. 3(b) and 3(e), this behavior likely stems from less accurate Stage-I angle estimates at low SNR, which limits the effectiveness of Stage-II refinement.
@@ V-C Performance with Different Range Parameters
We now investigate the impact of the distance between the scatterers and the BS on the performance of the proposed near-field channel estimation method. The distance varies from 5m to 120m, covering the near-field region and extending toward the far-field boundary. In the simulations, we set SNR = $30$ dB. Fig. 4(a) shows the NMSE of channel estimation versus distance. The NMSE of ANM(SI) remains nearly constant at approximately 25-25 dB across the entire range, indicating that the Stage-I reconstruction is largely insensitive to user location. In contrast, ANM(SI+SII) achieves substantially lower NMSE once the distance exceeds 20m, demonstrating the effectiveness of the Stage-II refinement when the user is not extremely close to the BS. The on-grid P-SOMP suffers from a persistent high error floor due to basis mismatch, whereas SW-OMP and SS-SIGW-OLS improve with increasing distance but remain ineffective in the deep near-field region.
@@ V-D Performance with Different Pilot Lengths
We next evaluate the impact of pilot length $Q$ on the performance of the proposed near-field channel estimation method. The pilot length varies from $Q=16$ to $Q=64$, corresponding to different training overheads. We fix the SNR at 30 dB. As shown in Fig. 5(a), the NMSEs of both ANM(SI) and ANM(SI+SII) improve steadily as $Q$ increases, and the curves gradually saturate for larger $Q$. These results demonstrate that the proposed methods effectively exploit additional pilot observations. In contrast, the on-grid P-SOMP exhibits only marginal improvement with increasing pilot lengths, whereas SW-OMP and SS-SIGW-OLS maintain nearly constant and significantly higher NMSE across all values of $Q$. The angle and range estimation results in Figs. 5(b) and 5(c) follow a similar trend. Both ANM(SI) and ANM(SI+SII) benefit substantially from longer pilot sequences, further confirming that the proposed algorithms capture the spherical wavefront structure more accurately as more observations become available.
@@ VI Conclusion
$0\in\partial\|\widehat{\mathbf{X}}\|_{\mathcal{A}}+\frac{1}{\tau}\mathcal{P}^{\ast}_{y}(\mathcal{P}_{y}(\widehat{\mathbf{X}})-\mathbf{y}).$ (-B.1)
There exists a dual variable $\mathbf{S}\in\partial\|\widehat{\mathbf{X}}\|_{\mathcal{A}}$ with $\|\mathbf{S}\|^{\ast}_{\mathcal{A}}\leq 1$ and $\langle\mathbf{S},\widehat{\mathbf{X}}\rangle_{\mathbb{R}}=\|\widehat{\mathbf{X}}\|_{\mathcal{A}}$. Thus, we have $\mathbf{S}=-\frac{1}{\tau}\mathcal{P}^{\ast}_{y}(\mathcal{P}_{y}(\widehat{\mathbf{X}})-\mathbf{y})$.
@@
If the condition $\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast}\leq\ \tau$ holds, inserting (-B.5) into (-B.4) leads to (24) in Theorem 1.
-C Proof of Theorem 2
We bound $\mathbb{E}\|\mathcal{P}_{y}^{\ast}(\mathbf{(}n)\|_{\mathcal{A}}^{\ast}$ by following a similar procedure in Appendix C and Appendix D of [Bhaskar-TSP13]. Notice that
$(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})^{2}$ $=\sup_{\omega\in[0,1)}\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\overline{\mathbf{a}(\omega)}\|_{2}^{2}$
$=\sup_{\omega\in[0,1)}\sum_{p=0}^{P-1}\left|\sum_{n=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}[\mathbf{q}_{n}]_{p}e^{-j2\pi\omega n}\right|^{2}$
$=\sup_{\omega\in[0,1)}\sum_{p=0}^{P-1}\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}\overline{[\mathbf{A}^{H}\mathbf{n}]_{m}}$
$~~~~~~~~~~~\cdot[\mathbf{q}_{n}]_{p}\overline{[\mathbf{q}_{m}]_{p}}e^{j2\pi\omega(m-n)}.$
Define
$W_{N}(e^{j2\pi\omega})$ $=\sum_{p=0}^{P-1}\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}\overline{[\mathbf{A}^{H}\mathbf{n}]_{m}}$ (-C.1)
$~~~~~~~\cdot[\mathbf{q}_{n}]_{p}\overline{[\mathbf{q}_{m}]_{p}}e^{j2\pi\omega(m-n)}.$
For any $\omega_{1},\omega_{2}\in[0,1)$, we have
$W_{N}(e^{j2\pi\omega_{1}})-W_{N}(e^{j2\pi\omega_{2}})$ (-C.2)
$\leq$ $|e^{j2\pi\omega_{1}}-e^{j2\pi\omega_{2}}|\sup_{\omega\in[0,1)}\left|W_{N}^{\prime}(e^{j2\pi\omega})\right|$
$=$ $2|\sin(\pi(\omega_{1}-\omega_{2}))|\sup_{\omega\in[0,1)}\left|W_{N}^{\prime}(e^{j2\pi\omega})\right|$
$\leq$ $2\pi|\omega_{1}-\omega_{2}|\sup_{\omega\in[0,1)}\left|W_{N}^{\prime}(e^{j2\pi\omega})\right|$
$\leq$ $4\pi N|\omega_{1}-\omega_{2}|\sup_{\omega\in[0,1)}W_{N}(e^{j2\pi\omega}),$
where the first inequality uses the mean value theorem and the last inequality follows from Bernstein’s theorem (see, for example, Theorem 3 in Appendix C of [Bhaskar-TSP13]) and the fact that $\underset{\omega\in[0,1)}{\sup}W_{N}(e^{j2\pi\omega})$ from (-C.1) is nonnegative. Defining the grid values $0,\frac{1}{T},\cdots,\frac{T-1}{T}$ and using (-C.2), it yields
$\sup_{\omega\in[0,1)}W_{N}(e^{j2\pi\omega})\leq\max_{t=0,\cdots,T-1}W_{N}(e^{j2\pi\frac{t}{T}})$ (-C.3)
$~~~~~~~~~~~~~~~~~~~~~~~+\frac{2\pi N}{T}\sup_{\omega\in[0,1)}W_{N}(e^{j2\pi\omega}),$
which leads to
$\sup_{\omega\in[0,1)}W_{N}(e^{j2\pi\omega})$ $\leq\left(1-\frac{2\pi N}{T}\right)^{-1}\max_{t=0,\cdots,T-1}W_{N}(e^{j2\pi\frac{t}{T}})$ (-C.4)
$\leq\left(1+\frac{4\pi N}{T}\right)\max_{t=0,\cdots,T-1}W_{N}(e^{j2\pi\frac{t}{T}}),$
where the last inequality holds when $T\geq 4\pi N$. Since $(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})^{2}=\underset{\omega\in[0,1)}{\sup}W_{N}(e^{j2\pi\omega})$, we bound $\mathbb{E}(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})^{2}$ by upper bounding $\mathbb{E}\underset{t=0,\cdots,T-1}{\max}W_{N}(e^{j2\pi\frac{t}{T}})$, where the expectation is with respect to the noise $\mathbf{n}$. Note that
$\mathbb{E}\max_{t=0,\cdots,T-1}W_{N}(e^{j2\pi\frac{t}{T}})$ (-C.5)
$=$ $\mathbb{E}\max_{t=0,\cdots,T-1}\sum_{p=0}^{P-1}\left|\sum_{n=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}[\mathbf{q}_{n}]_{p}e^{-j2\pi\frac{t}{T}n}\right|^{2}$
$\leq$ $\sum_{p=0}^{P-1}\mathbb{E}\max_{t=0,\cdots,T-1}\left|\sum_{n=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}[\mathbf{q}_{n}]_{p}e^{-j2\pi\frac{t}{T}n}\right|^{2}.$
Define $x_{p,t}=\sum_{n=0}^{N-1}[\mathbf{A}^{H}\mathbf{n}]_{n}[\mathbf{q}_{n}]_{p}e^{-j2\pi\frac{t}{T}n}$. It is easy to see $\mathbb{E}(x_{p,t})=0$. Denote $\mathbf{\widetilde{q}}_{n}]_{p,t}=[\mathbf{q}_{n}]_{p}e^{-j2\pi\frac{t}{T}n}$. For notational convenience, we also define $\mathbf{a^{\prime}}_{m}^{T}$ as the $m$th row of $\mathbf{A}$ and $\mathbf{g^{\prime}}_{p,t}$ a length-$N$ column vector with its $n$th entry being $\mathbf{\widetilde{q}}_{n}]_{p,t}$. Then
$\text{Var}(x_{p,t})$ $=\text{Var}\left(\sum_{m=0}^{M-1}\sum_{n=0}^{N-1}\overline{[\mathbf{A}]}_{m,n}[\mathbf{\widetilde{q}}_{n}]_{p}[\mathbf{n}]_{m}\right)$ (-C.6)
$=\text{Var}\left(\sum_{m=0}^{M-1}\mathbf{a^{\prime}}^{H}_{m}\mathbf{g^{\prime}}_{p,t}[\mathbf{n}]_{m}\right)$
$=\sum_{m=0}^{M-1}|\mathbf{a^{\prime}}^{H}_{m}\mathbf{g^{\prime}}_{p,t}|^{2}\sigma^{2}=\|\overline{\mathbf{A}}\mathbf{g^{\prime}}_{p,t}\|_{2}^{2}\sigma^{2}$
Denote $\sigma_{p,t}^{2}=\|\overline{\mathbf{A}}\mathbf{g^{\prime}}_{p,t}\|_{2}^{2}\sigma^{2}$ and $\sigma_{p}^{2}=\|\mathbf{A}\|_{2}^{2}\|\mathbf{g}_{p}\|_{2}^{2}\sigma^{2}$, where $\mathbf{g}_{p}$ is the $p$th column of $\mathbf{G}$. Thus, $x_{p,t}\sim\mathcal{CN}(0,\sigma_{p,t}^{2})$ and $\frac{2|x_{p,t}|^{2}}{\sigma_{p,t}^{2}}$ is a chi-squared random variable with two degrees of freedom. Furthermore, we have
$\mathbb{E}\left[\max_{0,\cdots,T-1}2|x_{p,t}|^{2}\right]=\int_{0}^{\infty}\mathbb{P}\left\{\max_{t=0,\cdots,T-1}2|x_{p,t}|^{2}\geq t\right\}dt$ (-C.7)
$=$ $\int_{0}^{\delta_{p}}\mathbb{P}\left\{\max_{t=0,\cdots,T-1}2|x_{p,t}|^{2}\geq t\right\}dt$
$~~~~~~~~~~~~~~~~~+\int_{\delta_{p}}^{\infty}\mathbb{P}\left\{\max_{t=0,\cdots,T-1}2|x_{p,t}|^{2}\geq t\right\}dt$
$\leq$ $\delta_{p}+T\int_{\delta_{p}}^{\infty}\mathbb{P}\left\{\frac{2|x_{p,t}|^{2}}{\sigma_{p,t}^{2}}\geq\frac{t}{\sigma_{p,t}^{2}}\right\}dt$
$\leq$ $\delta_{p}+T\int_{\delta_{p}}^{\infty}\mathbb{P}\left\{\frac{2|x_{p,t}|^{2}}{\sigma_{p,t}^{2}}\geq\frac{t}{\sigma_{p}^{2}}\right\}dt$
$=$ $\delta_{p}+T\int_{\delta_{p}}^{\infty}e^{-\frac{t}{2\sigma_{p}^{2}}}dt=\delta_{p}+2T\sigma_{p}^{2}e^{-\frac{\delta_{p}}{2\sigma_{p}^{2}}},$
where the second inequality above uses the fact that $\sigma_{p,t}^{2}\leq\sigma_{p}^{2}$. By choosing $\delta_{p}=2\sigma_{p}^{2}\log T$ and $T=4\pi N\log N$, we have
$\mathbb{E}(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})$
$\leq$ $\sqrt{1+\frac{1}{\log N}}\sqrt{\left(\frac{1}{2}\sum_{p=0}^{P-1}\mathbb{E}\left[\max_{0,\cdots,T-1}2|x_{p,t}|^{2}\right]\right)}$
$=$ $\sqrt{1+\frac{1}{\log N}}\sqrt{\sum_{p=0}^{P-1}\left(\sigma_{p}^{2}\log(4\pi N\log N)+\sigma_{p}^{2}\right)}$
$=$ $\sqrt{1+\frac{1}{\log N}}\sqrt{\log N+\log(4\pi\log N)+1}\sqrt{\sum_{p=0}^{P-1}\sigma_{p}^{2}}$
$=$ $\sqrt{1+\frac{1}{\log N}}\sqrt{\log N+\log(4\pi\log N)+1}\|\mathbf{A}\|_{2}\|\mathbf{G}\|_{F}\sigma$
where the first inequality above follows directly from (-C.4), (-C.5) and (-C.7). This completes the proof. $\square$